OpenStudy (anonymous):

Fool's Riddle of the day, Anthony was travelling to the city of Far Away. Along the way he met two other travelers, Basil and Cecil, who were headed in the same direction. When they stopped for the night, Anthony realised that he had brough no food. Basil had packed four sandwiches while Cecil had packed three. They agreed that all there of them would share the food equally, and Anthony would pay 7 silve coins as his share. How would Basil and Cecil divide the money?

OpenStudy (kunal):

basil will get 4 and cecil will get 3 ........... is it the answer??

OpenStudy (anonymous):

Basil will get 5 and Ceil will get 2

OpenStudy (anonymous):

Because each will get 7/3 food....... For Anthony's food ...... Basill will give (4-7/3)=5/3 food and Ceil will give (3-7/3)=2/3 food....